OpenStudy (anonymous):

What must be added to 3x -7 to make x^2 + 4x -1?

OpenStudy (anonymous):

@amistre64

rishavraj (rishavraj):

3x - 7 + (y) = x^2 + 4x - 1 y = x^2 + 4x - 1 - 3x + 7

rishavraj (rishavraj):

so it would be x^2 + x + 6

OpenStudy (anonymous):

What about this one? By how much does a^3 - 2a exceed a^2 + a - 6?

rishavraj (rishavraj):

it would be a^3 - 2a - (a^2 + a - 6) = a^3 - 2a -a^2 - a + 6 = a^3 - a^2 - 3a + 6

OpenStudy (anonymous):

Which one is it?

rishavraj (rishavraj):

the answer a^3 - a^2 - 3a + 6

OpenStudy (anonymous):

The perimeter of a triangle is 7a - 11b. If two of its sides are 2a + b and a - 9b, what is the third side?

rishavraj (rishavraj):

2a + b + a - 9b + (y) = 7a - 11b y = 7a - 11b - 3a + 8b y = 4a - 3b